# Same model, better performance: the impact of shuffling on DNA Language Models benchmarking

### v1

@@ 2.3 Our contributions
Finally, we evaluate the impact of pre-shuffling the CpG methylation task data on two additional LM architectures, DNABERT-2 [4] and ResNet-LM, a baseline model proposed by [1]. Code is publicly available at https://github.com/baillielab/BEND.
@@ 3.2 CpG methylation task greatly benefits from shuffling
The reasons for this increase in performance could be the high autocorrelation in the DNA sequences of the CpG methylation task. For each task, we computed how many consecutive sequences overlap and the median percentage of overlapping length (see Appendix Table 3). In the CpG methylation task 51.9% of consecutive sequences overlap by at least one nucleotide. The median of the overlapping length is of 449 nucleotides, equal to 87.7% of the entire sequence length.
@@ 4 Conclusion
Finally, we have shown that pre-shuffling the data is a simple fix for disentangling benchmark performance from hardware-specific hyperparameters. We hope this paper provides a practical example highlighting the difficulty of properly implementing benchmarking frameworks, the need for apprechiating domain specific knowledge and best practice, and more broadly the importance of following best practices and the unintended consequences when they are not respected.
@@ 5 Acknowledgments
This work was carried out with funding by the Baillie Gifford Pandemic Science Hub. We would like to thank our colleagues for the invaluable discussions and feedback that contributed to the creation of this paper.
@@ Appendix A Appendix
Task Pct. of overlapping sequences Median pct. of shared nucleotides Weighted pct. of overlapping nucleotides
Histone modification 17.03 19.92 3.39
Gene finding 7.09 12.39 0.88
Enhancer annotation 1.75 49.27 0.86
CpG methylation 51.88 87.70 45.50
Chromatin accessibility 28.29 20.31 5.75

### v2

@@ 2.3 Our contributions
Finally, we evaluate the impact of pre-shuffling the CpG methylation task data on two additional LM architectures, DNABERT-2 [4] and ResNet-LM, a baseline model proposed by [1].
Code is publicly available at https://github.com/baillielab/BEND.
@@ 3.2 CpG methylation task greatly benefits from shuffling
The reasons for this increase in performance could be the high autocorrelation in the DNA sequences of the CpG methylation task. For each task, we computed how many consecutive sequences overlap and the median percentage of overlapping length (see Appendix Table 3). In the CpG methylation task 37.3% of consecutive sequences overlap by at least one nucleotide. The median of the overlapping length is of 462 nucleotides, equal to 90.3% of the entire sequence length.
@@ 4 Conclusion
Finally, we have shown that pre-shuffling the data is a simple fix for disentangling benchmark performance from hardware-specific hyperparameters. We hope this paper provides a practical example highlighting the difficulty of properly implementing benchmarking frameworks, the need for appreciating domain specific knowledge and best practice, and more broadly the importance of following best practices and the unintended consequences when they are not respected.
@@ 5 Acknowledgments
This work was carried out with funding by the Baillie Gifford Pandemic Science Hub. This work was supported by the Edinburgh International Data Facility (EIDF) and the Data-Driven Innovation Programme at the University of Edinburgh. Access to EIDF was facilitated through the University of Edinburgh’s Generative AI Laboratory GAIL Fellow scheme. We would like to thank our colleagues for the invaluable discussions and feedback that contributed to the creation of this paper.
@@ Appendix A Appendix
Task Pct. of overlapping sequences Median pct. of overlapping nucleotides Weighted pct. of overlapping nucleotides
Histone modification 16.81 19.73 3.32
Gene finding 2.99 9.83 0.29
Enhancer annotation 2.81 54.44 1.53
CpG methylation 37.34 90.23 33.70
Chromatin accessibility 28.11 20.12 5.66